OpenStudy (anonymous):

z=1-cosθ+isinθ,then |z|=

ganeshie8 (ganeshie8):

z = a+ib |z| = sqrt(a^2 + b^2)

ganeshie8 (ganeshie8):

z=1-cosθ+isinθ a = ? b = ?

OpenStudy (anonymous):

then

ganeshie8 (ganeshie8):

z=1-cosθ+isinθ notice that `1-cosθ` is not attached to `i`, so it becomes the real part

ganeshie8 (ganeshie8):

a = 1-cosθ b = sinθ

ganeshie8 (ganeshie8):

|z| = sqrt(a^2 + b^2) = ?